OpenStudy (anonymous):

since e^0=1 what is e^1=

OpenStudy (dumbcow):

just e^1 or 2.72

OpenStudy (anonymous):

actual am solving a calculus problem) of derivative

OpenStudy (anonymous):

solving for y,inother to get the equation of a tangent line

OpenStudy (anonymous):

y=x^3 e^x when x=1

OpenStudy (anonymous):

if i substitute the 1 into x my answer comes out as "e"

OpenStudy (dumbcow):

ok so y' wil give you the slope y' = 3x^2*e^x + x^3*e^x x=1 y'(1) = 3e +e = 4e

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

just wanted to be sure if i were doing the right thing

OpenStudy (dumbcow):

no problem :)